OpenStudy (anonymous):

How is the natural logarithm generally written? Explain why the natural log of e is equal to 1.

OpenStudy (anonymous):

ln(x)

OpenStudy (anonymous):

for the first ? right?

OpenStudy (zzr0ck3r):

ln(x) ln(e) = 1 e^(?) = e ? = 1

OpenStudy (anonymous):

1

OpenStudy (zzr0ck3r):

thats the explination of why

OpenStudy (anonymous):

great, you are so helpful!

OpenStudy (zzr0ck3r):

log_a(x) = y is the same thing as a^y = x so log_e (e) = 1 is e^1 = 1

Parth (parthkohli):

\( \color{Black}{\Rightarrow \ln e =\log_ee}\) Anything to the power of 1 produces the same number! :)

OpenStudy (zzr0ck3r):

sorry that last line should be e^1 = e

OpenStudy (anonymous):

ln(e) = 1 e^(1) = e 1= 1 is this a good explanation?

OpenStudy (anonymous):

or should i go with the one from @ParthKohli?

OpenStudy (zzr0ck3r):

log_a(x) = y is the same thing as a^y = x so log_e (e) = 1 is e^1 = e I would use something more along the lines of this, I think understanding his takes a little more comfort with logs.

OpenStudy (zzr0ck3r):

the first three lines should say it all

OpenStudy (zzr0ck3r):

log_a is log base a btw

OpenStudy (zzr0ck3r):

well it is for that explination:)

OpenStudy (anonymous):

thank you so much!!!!